OpenStudy (anonymous):

what is the equation of the line that crosses trough the points (3,0) and (-3,6)?

OpenStudy (anonymous):

a.Y= -x +3 b. Y= x + 3 c. Y= -x d. Y= -3x ... which of these.

OpenStudy (anonymous):

Grandient is the change in y over the change in x. The change in y here is -6 (0-6=-6) and the change in x is 6 (3-(-3)=6) so it's -6/6 which is -1. That means it has -x in it. So you know the y value goes down by one every time the x value goes up 1. So is you take (-3,6), you need to move 3 along the x to get to 0 so the y value goes down by 3 to get 3. So the gradient is -1 and the intercept is 3. So, y= -x + 3

OpenStudy (anonymous):

thanks so much